NMR Blog

Bruker 800 NMR Cryoprobe Test Results

From Dr. Seho Kim:



All targeted pw90’s can be achieved within the allowed power levels. Ubiquitin spectra are OK too. I have no other issue at this moment except for the lineshape related to the temperature gradient in the probe. The cryoprobe is ready to use.

Items (specifications): measured values

CHCl3 linshape (0.8/8.0/16.0 Hz): 1.08/7.4/16.9 Hz

CHCl3 lineshape with spin (0.7/7.0/14.0 Hz): 0.64/6.1/14.3 Hz

Sucrose sensitivity (1200): 1212 (Bruker measured 1280)

Water suppression (80/160 Hz): 33.7/88.3 Hz

1H pw90 (8.0u at <15W): 7.5u at 10W

13C pw90 (12.0u at <205W): 11.4u at 150W

15N pw90 (32.0u at <287W): 32u at 270W

2H pw90 (100.0u at <28W): 88.7u at 24W

Gradient strength(>0.66 T/m): 0.67 T/m

Gradient recovery (>95%, <100u): >95%, <10u

Overnight stability (>95%): 96.5%

Magnet drift:0.8 Hz/hr

Ubiquitin spectra: presat1D, 15N-HSQC, 13C-HSQC, HNCO, HN(CA)CO, HNCA, HN(CA)CO,


               TOCSY-15N-HSQC, NOESY-15N-HSQC


800 NMR With Bruker Console

The 800 NMR is working with the new Bruker console. The system currently has a room temperature probe and soon the cryoprobe will be installed.

Bruker 600V Specs - posted on OCTOBER 2016

Bruker 600V – Probe Z816801_0140 PA TXI 600S3 H-C/N-D-05-Z
Lineshape 0.41/3.7/7.2 Hz (spin 20Hz)
Lineshape 0.43/3.6/7.1 Hz (spin 0)
S/N EtB (2ppm) 1187.6:1
S/N EtB (200Hz) 1307:1
Sucrose S/N 335.2:1
P90 1H 7.8us @ 13.2W
P90 13C 11.8us @ 170W
P90 15N 36.3us @ 164W

Varian 600 became Bruker AVANCE III HD 600

As of October 13, 2016 the Varian 600 VNMRS console was replaced with Bruker AVANCE III HD console. The system is working good.

Bruker 600B Specs - March 1, 2018


600B specifications: All specs are good.

CHCl3 lineshape (0.65 / 6.0 /12.0 Hz); 0.61 / 4.8 / 9.3

Sucrose sensitivity (>275); 318.7

Magnet Drift: 0.04 Hz/h

System stability (>95%): >99%

1H pw90: 8.0 at -12.4

13C pw90: 11.3 at -23.0

15N pw90: 32.5 at -24.4

2H pw90: 78.5 at -14.7

Bruker 600B Specs - August 2017

Basic specs of 600B were tested and results are as good as stated in acceptance tests.
CHCl3 lineshape (0.6 / 6.0 /12.0 Hz); 0.48/5.3/11.4
Sucrose sensitivity (>275); 320.8
Magnet Drift: 0.13 Hz/h
System stability (>95%): >99%
1H pw90: 8.0 at -12.4
13C pw90: 11.4 at -23.0
15N pw90: 33.2 at -24.2
2H pw90: 80 at -14.7

Bruker 600 - Tests Feb 2017

These are the specifications data of 600B taken during the week of Feb 6 and all looked OK.
CHCl3 line shape (0.65 / 6 /12 Hz); 0.6/4.7/9.2
Sucrose sensitivity (SNR=275); 302.9
Magnet Drift; 0.3 Hz/hr
Overnight Stability; 99%
H1 pw90; 8.0 us at -12.4 dB
C13 pw90; 11.3 us at -23.0 dB
N15 pw90; 32.7 us at -24.4 dB
H2 pw90; 80 us at -14.7 dB

Seho Kim

Bruker 600 Software update

The 600B spectrometer now has Topspin 3.5 PL6.

Specs-tests 600B Feb 2016

Bruker 600 NMR (by Seho Kim)
Drift; < 1Hz/hr;
Stability (>99%); >99.2
Lineshape (0.6/6/12); 0.5 / 4.7 / 9.2
Sensitivity (275, sucrose); 299
H1 pw90; 8.1 at -12.4
N15 pw90; 33.0 at -24.4
C13 pw90; 13.3 at -23
H2 pw90; 80 at -14.7
gNhsqc; OK
gChsqc; OK

600B Field drift checked and fixed

The Bruker 600 field seemed to drift upfield at about 18Hz/hr. The background drift correction was set to -61 field units per hour. Resetting it to zero (or just -1) corrected the issue. It was reset by doing a 'ha' and then going into BSMS control and lock and then field control panel and editing the appropriate field value.

600B Software update

We have updated the 600B with new computer and Topspin 3.2 patch level 6. All the features of the system are working.

600B - backup

The directory /opt/topspin2.1.1/data on 600B was deleted for more disk space. Proteomics IT staff managed to copy it to /u1/home/nmrdata/prot-600b.

600B probe back

The Bruker 600 probe is back on Friday Nov.22. It worked fine over the weekend after some initial tuning issue.

RT Probe 1H not Tuning - Bruker 600

The proton tuning of the RT probe at Bruker 600 is not working. We checked the system and it is either the motor is stuck or the probe circuitry is bad or both. The console electronics/software seem fine as seen by tests with an external "probe circuit tunable to 600 MHz".

We have opened a ticket and the probe will be sent out to Bruker on 10/29/2013.

B-600 Software upgrade

The following is the message from Seho on 8/14/2013.

"Today I tried to upgrade Bruker 600 NMR with the new computer and Topspin 3.1.6 but I got the same problem before; The BSMS display did not show the main page. As recommended by Bruker, I attempted to upgrade the ELCB (BSMS CPU) but upgrading firmwares was not working. I installed Topspin 3.2.2 to see any change but ended to the same failure in the BSMS display. We may need a fresh installation of OS and Topspin or assistance from Bruker which costs about $2000.

Bruker 700 - Tests March 2017

These are measured specs. The sucrose spec is a little less than the original outstanding spec. (which is above the Bruker nominal spec.).
CHCl3 Lineshape (-/8/16); 1.08/6.4/12.8
Sucrose Sensitivity (1150); 1112
1H pw90; 7.8 at -12.04
13C pw90; 12.4 at -22.36
15N pw90; 26.5 at -22.89
2H pw90; 90.0 at -15.0
drift; 0.245 Hz/hr
stability; 96.1% with Z1 autoshim

Seho Kim

Bruker 700 Software update

The Bruker 700 now has Topspin 3.5 PL6.

Bruker 700 - IPSO reboot issue

The Bruker 700 IPSO unit wouldn't reboot. It got stuck at code C0 or d2. Resetting the firewall status at the host computer cleared the error.

700B - CryoProbe service done

The Bruker 700 cryoprobe service started on Monday May 16, 2016 and completed on Tuesday, May 17, 2016 afternoon.

Specs-tests 700B Feb 2016

Bruker 700 NMR (by Seho Kim)
Two minor issues (marked with * and **) will be checked again.
Drift; 0.89 Hz/hr
Stability (>98%); > 99%
Lineshape( 1*/8/16); 1.4*/7.0 /14.8; The original spec shows no value for linewidth at 50% (*)
Sensitivity (1150, sucrose); 946**; This is no change since the last year.
1H pw90; 7.8 at -12.04
13C pw90; 12.3 at -22.36
15N pw90; 26.5 at -22.89
2H pw90; 93 at -15.0
hsqcetfpf3gpsi; OK
hsqcetgp; OK

700B - Shared Cryo Probe Compressor swapped

The 700B outside compressor for the cryoprobe was swapped by Bruker.

700B - Shared Cryoprobe Annual Maintenance Update

The cryoprpbe was cooled down on Friday May 22, 2015. This time the cooldown/service took longer than usual as a filter pack had burst in the He supply line. The Cryobay was cleaned and the filterpack was removed from the supply line and reccoled. The coil heater is at 30% and preamp heater at 170%. The pulse widths have shortened by about 5% after service.

700B - Shared - Cryoprobe Annual Maintenance

The annual cryoprpobe maintenance was started on May 18,2015.

700B - Shared Shigemi Topshim 2

Bruker's Clemens Anklin evaluated the topshim data and gave the following response.

"the profiles look fundamentally ok but some are shifted.

Sucrose looks perfect.

700B - Shared Shigemi Topshim

Paolo brought to our attention an issue of cold probe shimming using topshim with our shared 700 (Bio NMR lab). The sample is protein in 90% H2O about 300ul in Bruker Shigemi tube. When we center the shigemi tube with the depth gauge coil center marking topshim B0 standard deviation after shimming is anywhere between 5 to 10 Hz. Seho and I could reproduce this observation.

700B - Shared: Cryo Probe Service

The annual cryo probe service was completed on Friday June20, 2014. The system is working fine.

700 - shared cryoprobe power is recycled - power failure

On March 26 there was a brief power outage and the power the shared 700 wasn't restored before the UPS quit. There was a fuse in the third floor panel room had tripped and that impacted the power for just this unit and the 800 magnet pump. After resetting the fuse power was restored and the cryoprobe warm-up was completed. The turbo pump was kept running overnight to achieve cold box vacuum and on March 27 cool-down is underway.

700-Shared - diagnostics - Seho's results

As per Seho's mail:


Data run during the weekend were stored under /data/Skim/Jerry101513;
20 trosyf3gpphsi19.2
21 trhncacbgp2h3d.x
22 trhncagp2h3d.x

700-Shared-diagnostics summary

A manostat was added to the 700B magnet on 10/14/2013. It did not alter the lock level drop and Z1 drift behavior. Opened a ticket (#150200) with Bruker on 10/16/2013. They analyzed the stability data and looked at stability data with just supercon shim and supercon shim + RT Z1 shim alone. Bruker's final conclusion is that the supercon Z1 needs touching. Currently, awaiting Bruker's recommendation for the course of action.

700-Shared-diagnostics -2

We set up again another stability test (10/9/2013) with Z1 auto shim and the results are encouraging. The lineshape and intensities were very stable that too in the initial part where we saw degradation before without Z1 shim. The lock level also remained steady.

Next course of action: In order to isolate if it is really Z1 shim or magnet pressure instability, we will install the manostat and repeat the stability test. This might help us narrow down the issue further.

700 - shared NMR diagnostics

We started (10/8/2013) the shim stability test using autotest sample that has 1% water, methanol, and acetonitrile in D2O. We started the experiment at 5:30PM and ran it for 12 hours. At the start, we had just closed the big outside door to the lab and the room temperature was, perhaps, still stabilizing. The 12 hour run results show a slow predominant Z1 shim change which was also pronounced in the earlier part of the evening. The line shape remained intact throughout except for the broadening effect. Up on correcting Z1 shim by about -90 units the shim level and lock level comes back.

Cryoprobe - NMR coil heater power - shared 700

There is a clear correlation between increased NMR bore temperature and a concomitant decrease in NMR coil heater power. The bore temperature seem to increase when higher sample temperature is set say above 25C. The preamp heater power is quite stable.

bore temperature/NMR coil heater/preamp power
301.86/17%/146% sample temp 25C
304.89/15%/145% sample temp 28C
305.04/16%/148% unknown sample temp (must be close to 28 C)
309.13/14%/147% sample temp 32C
327.34/7%/146% unknown sample temp

The temperature gas flow is 670 L/h.

Cryoprobe (readings unstable) - shared 700

The cryoprpbe coil heater power is dropping. After the re-installation 7/19/2013 the coil heater power was 21% and then it dropped to 19% on 7/22/2013 and on 7/23 it is at 7%. Bruker has been contacted.

Bruker response is that the readings are nothing to be alarmed and so we will wait and see.

Status update - shared 700

700B with the cryoprobe looks OK to use.
Filling helium will be scheduled on 7/25(Thu) for 600V (9am-11am), 600B (10-12), and 700B(1pm – 4pm).


Status update - shared 700

The Bruker service was done for replacing the He compressor and the 1H pw90 of the cryoprobe was tested quickly. The 1H pw90 was 8.24us, which is the same as the value we have when the cryoprobe runs normally. It looks replacing the He compressor fixed the problem. Tomorrow, I will continue some tests further and tomorrow afternoon 700B will be open to users.

Seho Kim

Status update - shared 700

Bruker engineer is working on the 700B cryoprobe system. The helium compressor was replaced in this morning and testing the cryoprobe will be followed sometime. 700B will be down until announcing the availability.

Seho Kim

Tuning issue (RT Probe) - shared 700

The 1H preamp and filters came in yesterday and these were installed. 700B will be ready to use today afternoon after checking pw, installing trosy relaxation experiments, and filling LN2. While Bruker engineers were checking the 700B2 cryoprobe, they helped to install the preamp and found that the tuning display by ATMM was not shown correctly when sw was reduced comparing to the default sw=15MHz. They think that it might be related to the software upgrade for the old HTTP/2 but tuning function and display are OK with default sw.

Tuning issue (RT Probe) - shared 700

Bruker thinks that the Proton preamp at the shared 700 is not working properly. Replacing that might solve the RT probe proton tuning issues. Awaiting a quote for the preamp.

CryoPlatform Service update - shared 700

From Seho:

Bruker recommends to replace the He compressor since the current service could not make the cryoprobe working normally. They think that the outdoor He compressor could not build up the right pressure. They will send another quote costing about $25,000.

In the meanwhile they are working here, they will diagnose the problem related to the RT probe.

RT Probe issues at shared 700

Yesterday (6/25/2013)– the Bruker guys did not have time to do the testing. They did it this afternoon. They tested the probe with Babis 700 console outside the magnet and there was no problem. We tested the probe outside the magnet with the shared 700 console again to eliminate any magnetic material in the proton coil that might have disturbed the tuning while the probe was inside the magnet and the tuning did not work. So the source of trouble seem to be with the console.

700 - Shared NMR is up

The shared 700 NMR is up and running with RT probe.

RT Probe Status - 700 Shared

Message from Seho Kim: The repaired RT probe looks OK to tune the 1H of salty samples. I will shim and check specs further. The 700B may be available tomorrow morning. Since the 700B topspin was upgraded to 3.2, please use the right version.

RT probe - 700 Shared

RT probe is back from Bruker. We are checking with Bruker about cryo platform service schedule to decide whether to install the RT probe at this time.

RT probe 700 - Shared

The RT probe is with Bruker to check tuning issues with salty samples.

Cryo platform service 700-Shared

Bruker has been contacted to do cryo platform service. They are in the process of scheduling the service and checking cryo probe performance.